# Adaptive behavior with stable synapses

### v2

@@ 3 Results
Figure 1: Dynamical adaptation of temporal trajectories: A. Overview of the network architecture employed. A recurrent network composed of $N=20$ units (described by the vector $\bm{z}(t)$, depicted in pink, which follows Eq. (6)) is utilized for learning a periodic trajectory, following the prescription of reservoir computing. One recurrent network is tasked with estimating the gradient of virtual weights, while another is dedicated to estimating the behavioral reconfiguration (illustrated in green and orange, respectively) resulting from these updated virtual weights. B. Illustration of our architecture dynamically adjusting (without synaptic alterations) to adhere to the desired dynamics. Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights $\delta w$, a $N$-dimensional vector (shown in the green trajectory). Subsequently, these updates are transmitted to the second network, modifying the decoding of reservoir dynamics (indicated by the orange lines). Initially, the reservoir receives the target trajectory as input (in open loop, before the red vertical line), which is later replaced by the estimated trajectory itself (in closed loop, after the red vertical line). C. Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (solid: median, dashed: 20-th/80-th percentile (statistics evaluated over 10 realizations).
@@
We consider the task of autonomously predicting a temporal trajectory $\{y^{targ}(t)\}_{t}$. Following the reservoir computing paradigm, we employ an RNN (Fig. 1A, pink), whose dynamics follows Equation (6), to extract temporal features. During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input ($x=y^{targ}(t)$ in Eq. (6), while $R$ is Gaussian matrix with zero mean and variance, $\sigma_{R}^{2}$) and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig. 1B). This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section, with the inputs $\bm{x}$ replaced by the features $\bm{z}(t)$ extracted by the RNN. The readout virtual weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$. This results in the following dynamics:
$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},\bm{w})\simeq{\color[rgb]{1,0.58984375,0.70703125}\bm{z}}\cdot\bm{w}\\ \tau_{w}\,\dot{\bm{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)\simeq{\color[rgb]{1,0.58984375,0.70703125}\bm{z}}\,e\end{cases}$ (9)
Here, the network $W_{\Theta^{w}}$$({\color[rgb]{1,0.58984375,0.70703125}\bm{z}},e)=\textrm{GMNN}_{\Theta^{w}}($$\bm{z}$$,e|\gamma_{e}e)$ is dedicated to estimating the gradient of virtual weights, while $Y_{\Theta^{y}}$(((($\bm{z}$$,w)=\textrm{GMNN}_{\Theta^{y}}($$\bm{z}$$,w|\gamma w)$ is tasked with estimating the predicted $y(t)$ as a function of the new virtual weights (Fig.1A green and orange respectively). Here, the parameters $\gamma,\gamma_{e}$ define the strength of the gain modulation, as explained in the Methods section.
@@
In the closed-loop phase, the features $\bm{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN ($x=y(t)$ in Eq. (6)). This results in an autonomous dynamical system that reproduces the target trajectory. (see Fig. 1B). In Fig.1B we report an example of successful dynamical adaptation of our model. Errors, represented by the blue trajectory, are fed back to the initial network to assess necessary updates to virtual weights, shown in the green trajectory. These updates are then transmitted to the second network, modifying the decoding of reservoir dynamics indicated by the orange lines. Initially, the reservoir receives the target trajectory as input in an open-loop fashion before the red vertical line, which is later replaced by the estimated trajectory itself in a closed-loop configuration after the red vertical line. The networks were pre-trained on five target frequencies marked by red vertical lines in Fig.1C and then tested across a range of frequencies. Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.1C, solid and dashed lines represent respectively, median and 20th/80th percentile range. Statistics is evaluated over 10 realizations of the experiment).
@@
3.1 Multi Armed Bandits
@@
In this simplified bandit scenario, where state information is absent, the virtual weights $\bm{w}$ directly parameterize the policy probabilities: $\bm{\pi}=\mathrm{softmax}{(\bm{w})}$. Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig. 2A, green). This serves as an ideal test bed for evaluating the network’s ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios.
To parameterize $W_{\Theta^{w}}$()($\cdot$)( ⋅ ) we employ a gain-modulated architecture $W_{\Theta^{w}}$ $(\bm{\pi},a,r)=\mathrm{GMNN}_{\Theta^{w}}(\bm{\pi},a,r|\gamma r)$, where $r$ is the reward and $a$ is the action. We train this network to approximate the policy gradient update rule. As training data, we use policy gradient estimates over one single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment. We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) models in out-of-distribution (OOD) environments. We train the network to approximate the policy gradient update rule using policy gradient estimates from a single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment. We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) in out-of-distribution (OOD) environments.
To systematically investigate the impact of gain modulation ($\gamma=1$) on out-of-distribution performance, we compare it with networks where the reward does not modulate the network gain ($\gamma=0$). For each case, we select the optimal hyperparameters through a grid-based search (additional details in the Appendix).
Our experimental findings are presented in Fig.2. First, comparing models of different sizes (number $N$ of hidden units), we find that models with gain modulation require significantly fewer neurons to achieve maximum scores in OOD environments (Fig.2C).
In Fig.2B, we report the regret per round distribution at the 100th round and compare it with the distribution obtained by policy gradient with the same learning rate and iterations. A gain-modulated network achieves a regret distribution comparable to the policy gradient in both ID and OOD environments, with a lower median. In contrast, networks without gain modulation show significantly higher regret. Analyzing the regret curves in Fig.2D, we observe that a model with gain modulation often learns a more data-efficient algorithm than its source, even in OOD environments. Conversely, a model without gain modulation fails to generalize to OOD environments and does not converge to zero regret.
@@
3.2 Reinforcement learning in a reaching task.
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [31]), set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility. The position of the agent is encoded separately using 25 input units $\bm{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
To accomplish this task we consider a network architecture composed of two networks. One GM-network, $W_{\Theta^{w}}$$(r,a,\bm{\pi})=\textrm{GMNN}_{\Theta^{w}}(r,a,\bm{\pi}|\gamma_{r}r)$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig. 3A, green network), while another GM-network $Y_{\Theta^{y}}$$(\bm{x},\bm{w})=\textrm{GMNN}_{\Theta^{y}}(\bm{x},\bm{w}|\gamma\bm{w})$ estimates policy reconfiguration (Fig. 3A, shown in orange) resulting from changes in virtual weights. Here, the parameters $\gamma,\gamma_{r}$ define the strength of the gain modulation, as explained in the Methods section. Their dynamics can be described by the following equation Eq.(5). Indeed, we compare the case with (Fig. 3A) and without (Fig. 3B) gain modulation. We refer to Supporting Information for further details on the training procedure.
@@
We demonstrate, that our architecture is capable to perform the temporal computation required to learn delayed action-reward temporal relations (see section C in Appendix), requiring evaluating temporal credit assignment.
4 Discussion
It is believed that the remarkable capability of transformers to adapt to contextual information, is the key ingredient that allowed transformer-based architectures [4, 32], achieving state-of-the-art in many domains, such as processing and generation of natural language and images [6, 7, 8]. Currently, mechanisms based on attention, associative memory [9], and induction-based copying [10] by heads are the predominant tentative explanations for the emergence of in-context learning in transformers. This explanation theoretically demonstrates that the attention mechanism has ability to execute gradient descent updates [11]. However, the emergence of this property remains not fully understood and it is not clear how to translate this to the in-context learning capabilities observed in human and animals. In this this work we proposed a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios.
@@ 4 Discussion
When tasked to learn temporal trajectories, through error feedback and virtual weight updates, the network achieves successful dynamical adaptation, without synaptic plasticity. Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, that are updated in function of the reward. We stress that virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other RNNs, that were pre-trained to perform gradient updates. Another pre-trained network receives those weights as an input, along with the current environment features, and computes an adapted in-context response. We find that networks with gain modulation exhibit improved performance and robustness to variations compared to those without gain modulation. Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing its capability to face novel scenarios.
@@
In conclusion, our approach provides an explicit framework to induce ICL in biologically plausible networks, possibly opening the route to a formal understanding of ICL in biological agents.
Limitations of this study
In this study, we focused on dynamical adaptation of virtual readout weights, in other words, only the last layer is considered. We are neglecting the dynamics on recurrent weights and weights of the hidden units, which, in general, strongly affect performance and stability. Generalizability to untested real-world conditions is uncertain, and we do not investigate the scalability to larger architectures. We discuss that the mechanism we discuss could be present in biological agents, allowing in-context learning. However, the theoretical basis for attention mechanisms executing gradient descent updates lacks comprehensive empirical validation in biological experiments.
Additionally, the study does not fully address how dynamic gradient structures could be learned or developed in biological networks.
5 Acknowledgments

### v4

@@ Results
Figure 2: Dynamical adaptation of temporal trajectories: A. Overview of the network architecture employed. A recurrent network composed of $N=20$ units (described by the vector $\mathbf{z}(t)$, depicted in pink, which follows Eq. (4)) is utilized for learning a periodic trajectory, following the prescription of reservoir computing. One recurrent network is tasked with estimating the gradient of virtual weights, while another is dedicated to estimating the behavioral reconfiguration (illustrated in green and orange, respectively) resulting from these updated virtual weights. B. Illustration of our architecture dynamically adjusting (without synaptic alterations) to adhere to the desired dynamics. Errors (represented by the blue trajectory) are fed back to the initial network to assess necessary updates to the virtual weights $\delta w$, a $N$-dimensional vector (shown in the green trajectory). Subsequently, these updates are transmitted to the second network, modifying the decoding of reservoir dynamics (indicated by the orange lines). Initially, the reservoir receives the target trajectory as input (in open loop, before the red vertical line), which is later replaced by the estimated trajectory itself (in closed loop, after the red vertical line). C. Our networks were pre-trained on five target frequencies (marked by red vertical lines) and tested across a range of frequencies, evaluating the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (solid: median, dashed: 20-th/80-th percentile (statistics evaluated over 10 realizations).
@@
We consider the task of autonomously predicting a temporal trajectory $\{f(t)=y^{targ}(t)\}_{t}$. Following the reservoir computing paradigm (see Methods section), we employ a random RNN (Fig. 2A, pink), whose dynamics follows Equation (4), to extract temporal features $\mathbf{z}(t)$. During training, the reservoir receives the target dynamics $y^{targ}(t)$ as input ($x(t)=y^{targ}(t)$ in Eq. (4)) and is tasked with predicting the subsequent step of the trajectory via a linear readout of its activity (open loop, see Fig. 2B). This setup can be reformulated as a dynamical supervised learning problem, as described in the methods section. The readout weights can then be dynamically adjusted, minimizing the error between the target and the current prediction $y(t)$. This results in the following dynamics:
$\begin{cases}e&=(y^{targ}-y)\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}},\mathbf{w}\})\simeq{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}}\cdot\mathbf{w}\\ \tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}},e\})\simeq{\color[rgb]{1,0.58984375,0.70703125}\mathbf{z}}\,e\end{cases}$ (2)
[…]
@@
In the closed-loop phase, the features $\mathbf{z}$ are obtained by directly feeding the network estimation $y(t)$ as the input to the RNN ($x=y(t)$ in Eq. (4)). This results in an autonomous dynamical system that reproduces the target trajectory. (see Fig. 2B). In Fig.2B we report an example of successful dynamical adaptation of our model. Errors, represented by the blue trajectory, are fed back to the initial network to assess necessary updates to virtual weights, shown in the green trajectory. These updates are then transmitted to the second network, modifying the decoding of reservoir dynamics indicated by the orange lines. Initially, the reservoir receives the target trajectory as input in an open-loop fashion before the red vertical line, which is later replaced by the estimated trajectory itself in a closed-loop configuration after the red vertical line. The networks were pre-trained on five target frequencies marked by red vertical lines in Fig.2C and then tested across a range of frequencies. Evaluation was performed by assessing the mean squared error (MSE) between the target and estimated trajectories in closed-loop scenarios (see Fig.2C, solid and dashed lines represent respectively, median and 20th/80th percentile range. Statistics is evaluated over 10 realizations of the experiment).
@@
Extension to Reinforcement learning and the role of Gain modulation
@@
In this simplified bandit scenario, where state information is absent, the virtual weights $\mathbf{w}$ directly parameterize the policy probabilities: $\mathbf{\pi}=\mathrm{softmax}{(\mathbf{w})}$. Consequently, our focus lies primarily on analyzing the behavior of the network $W_{\Theta^{w}}^{\gamma}$()($\cdot$)( ⋅ ) acting on the virtual parameters (Fig. 3A, green). This serves as an ideal test bed for evaluating the network’s ability to learn the policy gradient update rule and generalize it to out-of-distribution scenarios. The network $W_{\Theta^{w}}^{\gamma}$$(\mathbf{\pi},a,r)$ has a gain-modulated architecture, as described in the Methods section, and is a function of the reward $r$ and the action $a$. Here $\gamma$ indicates the dependence on the gain-modulation strength. We train the network to approximate the policy gradient update rule using policy gradient estimates from a single learning trajectory (1000 rounds, learning rate $lr=0.1$) in an in-distribution environment. We then test the distilled policy gradient networks with a higher learning rate ($lr=1.0$) in out-of-distribution (OOD) environments. To systematically investigate the impact of gain modulation ($\gamma=1$) on out-of-distribution performance, we compare it with networks where the reward does not modulate the network gain ($\gamma=0$). For each case, we select the optimal hyperparameters through a grid-based search (additional details in the Appendix).
Our experimental findings are presented in Fig.3. First, comparing models of different sizes (number $N$ of hidden units), we find that models with gain modulation require significantly fewer neurons to achieve maximum scores in OOD environments (Fig.3C). In Fig.3B, we report the regret per round distribution at the 100th round and compare it with the distribution obtained by policy gradient with the same learning rate and iterations. A gain-modulated network achieves a regret distribution comparable to the policy gradient in both ID and OOD environments, with a lower median. In contrast, networks without gain modulation show significantly higher regret. Analyzing the regret curves in Fig.3D, we observe that a model with gain modulation often learns a more data-efficient algorithm than its source, even in OOD environments. Conversely, a model without gain modulation fails to generalize to OOD environments and does not converge to zero regret.
@@
Reinforcement learning in a reaching task.
We examine a reaching task, known as the dark room task (a simple instance of the water maze task [39]),
set in a 2D maze within the domain $(-1,1)\times(-1,1)$ with a grid size of $0.1$. Within this grid-like environment, the agent navigates by selecting one of four actions: up, down, left, or right, thereby determining its subsequent position. The primary goal is to locate a concealed object within the maze, with the agent having sole awareness of its own position. Feedback is provided via rewards, where the agent receives a reward of $10$ if the distance to the object is $0.1$, $15$ if the distance is 00, and 00 otherwise. Through iterative exploration, the agent develops a strategy to efficiently traverse the maze and pinpoint the object despite the limited visibility. The position of the agent is encoded separately using 25 input units $\mathbf{x}$ each, employing Gaussian activation functions distributed on a 5x5 grid in the maze and with a width of $0.2$.
To accomplish this task we consider a network architecture composed of two networks. One GM-network, $W_{\Theta^{w}}^{\gamma}$($(\{\mathbf{z},r,a,\mathbf{\pi}\})$ is responsible for estimating the necessary update of virtual weights through policy gradient (Fig. 4A, green network), while another GM-network $Y_{\Theta^{y}}^{\Gamma}$($(\{\mathbf{z},\mathbf{w}\})$ estimates policy reconfiguration (Fig. 4A, shown in orange) resulting from changes in virtual weights. Here, the parameters $\gamma,\Gamma$ define the strength of the gain modulation, as explained in the Methods section. Their dynamics can be described by the following equation Eq.(10). Indeed, we compare the case with (Fig. 4A) and without (Fig. 4B) gain modulation. We refer to Supporting Information for further details on the training procedure.
@@
Temporal credit assignment and Future Reward
We demonstrate, that our architecture is capable to perform the temporal computation required to learn delayed action-reward temporal relations (see section Reinforcement Learning and the Discount Factor in Appendix), requiring evaluating temporal credit assignment. Indeed, in reinforcement Learning (RL) an agent learns to maximize cumulative rewards. The discount factor, usually denoted by $\gamma_{d}$ (where $0\leq\gamma_{d}\leq 1$), is crucial in RL as it determines the importance of future rewards. A higher $\gamma_{d}$ values future rewards more significantly, encouraging the agent to consider the long-term consequences of its actions. Conversely, a lower $\gamma_{d}$ makes the agent prioritize immediate rewards. The cumulative reward $R_{t}$ at time step $t$ is given by:
$R_{t}=\sum_{s\geq t}r_{s}\gamma^{(s-t)}_{d}$ (3)
To account for this, the term $\mathbf{e}$ in Eq.(10) should be replaced by its exponential temporal filtering, with a timescale $\tau_{e}=-\frac{1}{\log(\gamma_{d})}$ (see also Eq.(13) in the Appendix). In this way actions in the past, can be potentiated when delayed rewards are received.
If the discount factor is zero, the agent’s policy would change only when it is very close to the moment the reward is received. In a reaching task, the agent will only change its policy when it is very close to the reward, completely ignoring the need for long-term planning and making it unlikely to ever reach the goal from distant starting points (see Fig. A2.D-E). As a result, the policy points towards the target position only when nearby the target itself (see Fig. A2F), resulting in failure when the agent randomly moves in the wrong direction at the beginning of the task (see Fig. A2F, black line).
On the other hand, recurrent connections enable a network to maintain a memory of past states and actions, effectively allowing it to use information from previous time steps to inform current decisions, achieving optimal planning and decisions even when far from the target location (and hence, from the reward, see Fig. A2.A-B-C).
Discussion
The brain’s ability to flexibly and rapidly adapt to novel tasks remains only partially understood. Different studies have highlighted that this adaptability relies on mechanisms such as dynamic network reconfiguration and context-dependent modulation, allowing agents to respond flexibly to environmental changes. However, a significant gap remains in understanding how these processes map onto specific neural substrates.
While synaptic plasticity has been extensively studied as a mechanism underlying learning and memory [40], emerging evidence suggests that it alone may not account for the full spectrum of adaptive behavior. Abraham et al. critically evaluate the synaptic plasticity and memory (SPM) hypothesis, concluding that non-synaptic mechanisms, such as intrinsic plasticity, structural neural changes, epigenetic regulation, glial contributions, network reorganization, and protein turnover, likely contribute significantly to memory storage and retrieval, thus broadening the framework for understanding dynamic adaptation. This perspective questions long-standing assumptions about neural circuits, highlighting that we may still lack full understanding of many processes we consider well-established.
Inspired by the brain’s adaptability, artificial neural network architectures have been challenged to replicate similar capabilities. Fast synaptic plasticity, hypothesized to orchestrate flexible cognitive functions, inspired several artificial architectures, such as fast-weight networks [41, 42] and hyper-networks that predict main network weights [43]. However, these approaches often rely on biologically unrealistic mechanisms, such as requiring neurons to directly act on the synapses of a second network or implementing "inner loops" in network dynamics [44, 43]. More biologically plausible alternatives, like coupled neural-synaptic dynamics [45], come closer to natural systems but still depend on complex credit assignment mechanisms.
Transformers, known for their remarkable in-context learning (ICL) capabilities, excel at tasks requiring adaptability, such as natural language and image processing [14, 15, 16]. Mechanisms such as attention, associative memory [17], and induction-based copying by attention heads [18] have been proposed to explain ICL in transformers, with theoretical work showing that attention mechanisms can implement gradient updates [19]. However, the emergence of these capabilities remains poorly understood, particularly when comparing them to the in-context learning observed in biological systems.
We propose a constructive method to induce in-context learning in biologically plausible neural networks, in a broad variety of scenarios. We offer an alternative formulation in which dynamic adaptation is modulated by dendritic input segregation and amplification. Virtual (fast) weights are encoded by the activity of a second network that modulates the transfer function of neurons in the base network. As a result, virtual weights are not synaptic weights, but their values and updates are evaluated and encoded by the activity of other neurons, which were pre-trained to perform gradient updates. Another pre-trained network receives those weights signals as an input, along with the current environment features, and computes an adapted in-context response.
This mechanism overcomes the strict locality constraints of synaptic plasticity, which require all elements involved in evaluating weight updates to be present at the synaptic site. Consequently, biologically plausible synaptic updates cannot explicitly rely on information from other weights, a requirement often seen in backpropagation methods. In contrast, virtual weights are encoded by neuronal activity and can propagate across the neural circuit, bypassing these limitations. This allows virtual weight updates to depend on other virtual weights, enabling a more natural and precise evaluation of credit assignment.
@@ Discussion
Our work generalizes and provides a biologically plausible implementation for deep gated RNNs performing ICL [46], maintaining a similar number of trainable parameters. By dividing the architecture into two components and introducing an additional objective that forces one network to approximate the gradient of the other, we achieve robust dynamic adaptation in tasks requiring temporal learning. When tasked to learn temporal trajectories, through error feedback and virtual weight updates, our network achieves successful dynamical adaptation, without synaptic plasticity. Similarly, in reinforcement learning scenarios, the policy in response to the environment state is dynamically modulated by virtual weights, which are updated in function of the reward. We find that networks with gain modulation exhibit improved performance and robustness compared to those without gain modulation. Moreover, these gain-modulated networks demonstrate more data-efficient learning algorithms, outperforming counterparts in both in-distribution and out-of-distribution environments, showing their capability to face novel scenarios.
While we do not fully address how the static weights of the networks involved can be learned using biologically plausible learning rules, we demonstrate that these architectures naturally emerge by tuning the readout weights of random networks. These weights could be shaped through evolutionary processes, adapted over slow timescales via reinforcement learning, or optimized using local plasticity rules based on feedback alignment. We argue that this represents a step forward in biological plausibility compared to existing deep-learning models like transformers, gated RNNs, and hyper-networks, which rely on more complex and biologically implausible credit assignment mechanisms such as backpropagation through time to train synaptic weights.
@@
In conclusion, our approach provides an explicit framework to induce in-context learning (ICL) in biologically plausible networks, offering a constructive pathway toward understanding ICL in biological systems and bridging the gap between neuroscience and machine learning.
Methods
Gain modulated reservoir computing (GM-RC)
Reservoir Computing
In this approach, a random RNN is employed to extract features from a time-dependent signal $\mathbf{x}(t)$. The dynamics of the RNN describe the evolution of $N$ hidden units $\mathbf{z}(t)=(z_{1}(t),\cdots,z_{N}(t))$, governed by the following differential equation:
$\tau_{z}\dot{\mathbf{z}}=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}$ (4)
The value of each unit represents the activity of a population of neurons following the Wilson and Cowan formulation ([47]). Here, $J$ and $R$ represent fixed random matrices, representing the recurrent connections and the projection from inputs to hidden units, respectively. Subsequently, the features extracted by the network can be utilized to predict a target signal $\mathbf{y}^{targ}(t)$ by learning readout weights $\Theta$, such that $|\mathbf{y}^{targ}(t)-\Theta\mathbf{z}(t)|^{2}$ is minimized. The reservoir is then implementing a map
$\mathbf{y}(t)=\mathrm{RNN}_{\Theta}(\{\mathbf{x}(s)\}_{s\leq t})$ (5)
where $\Theta$ represents trainable parameters. In the rest of the article, we will write $\mathbf{y}=\mathrm{RNN}_{\Theta}(\{\mathbf{x}\})$, dropping the temporal dependencies. 111 For this mapping to be well-defined, we can assume the RNN to have the echo-state property [48].
When addressing input-output mappings without time dependencies, we consider a network operating in the $\tau_{z},J_{ij}\to 0$ limit. In this scenario, the network computes an instantaneous function of the input, represented as $\mathrm{NN}_{\Theta}(x)=\Theta\phi\left(R\mathbf{x}\right)$. Essentially, this is equivalent to considering a one-layer feed-forward network with random fixed input weights. This architecture is also referred to in the literature as the Extreme Learning Machine [49].
Gain modulated network architecture
Analysing Equations (8) and (10) we observe that a network must possess the capability to perform multiplications of its inputs to approximate a gradient descent update of virtual parameters. Building upon this insight, we introduce a gain-modulated reservoir network (GM-RC). This architecture draws inspiration from the morphology and function of pyramidal neurons in the cortex, which nonlinearly integrate inputs from basal and apical dendrites [50, 51]. Here, we consider an additional input source $\mathbf{x}^{\text{ap}}$ randomly projected into the apical dendrite of each neuron by the matrix $R^{\text{ap}}$. Consistent with experimental observations in L5 pyramidal neurons, we allow the apical inputs to modulate the gain of the activation function, thereby altering its slope. Consequently, the resulting RNN equation is formulated as:
$\tau_{z}\dot{\mathbf{z}}$ $=\phi\left(\left(\mathbf{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}\right)\odot(J\mathbf{z}+\beta\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}+R\mathbf{x})\right)-\mathbf{z}$ (6)
Where $\mathbf{b}^{\text{ap}}$ is a constant bias vector. The hyperparameters $\beta,\gamma$ modulate the effect of the gain modulation of the apical inputs. Specifically, when $\gamma=0$, we obtain a network in which $\mathbf{x}^{\text{ap}}$ does not affect the gain modulation. Similarly, as before, the expression $\mathbf{y}=\mathrm{GMRNN}^{\gamma}_{\Theta}(\{\mathbf{x}|\mathbf{x}^{\text{ap}}\})$ will denote the input $\to$ output mapping implemented by a gain-modulated reservoir. As before, removing time dependencies, we will have an instantaneous function $\mathbf{y}=\mathrm{GMNN}^{\gamma}_{\Theta}(\mathbf{x}|\mathbf{x}^{\text{ap}})=\Theta\phi\left(\left(\mathbf{b}^{\text{ap}}+\gamma\cdot R^{\text{ap}}\mathbf{x}^{\text{ap}}\right)\odot(R\mathbf{x})\right)$ of the inputs $\mathbf{x}^{\text{ap}}$ and $\mathbf{x}$. We explicitly maintain the dependence on $\gamma$ because, in our experiments, we use this parameter to regulate the gain modulation effect of the apical inputs on the network. The choice for the name $\mathbf{x}^{\text{ap}}$ is inspired by the current received in the apical dendrites of L5 pyramidal neurons, that are believed to carry contextual/high-level information [50].
Dynamical supervised learning for temporal trajectory
We consider the task of learning a target temporal trajectory. Each environment $\alpha$ is defined by a target trajectory $y_{\alpha}^{targ}(t)$, such that the input and the feedback $x_{\alpha}(t)=f_{\alpha}(t)=y_{\alpha}^{targ}(t)$ are both given by the target sequence. The network response is defined as the current estimation of $y^{targ}(t)$ via a linear readout $y(t)=\mathbf{w}\cdot\mathbf{z}(t)$, where $\mathbf{z}$ are the activity of a random RNN that operates as a reservoir computer extracting nonlinear features of the input sequence. The RNN dynamics is given by equation (4), with input $x$ given by the target sequence. The readout weights in turn dynamically adapt by following the delta rule, minimizing the reconstruction error.
The learning process can then be described by the following dynamical system:
$\begin{cases}e&=(y^{targ}-y)\\ y&=\mathbf{z}\cdot\mathbf{w}\\ \tau_{z}\dot{\mathbf{z}}&=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}\\ \tau_{w}\,\dot{\mathbf{w}}&=\mathbf{z}\,e\end{cases}$ (7)
where we removed the dependence on time $t$ for simplicity. The operations required are nonlinear, and usually are viewed as implemented by the plasticity rule and by the multiplication between presynaptic activity and synaptic weights. However, here we interpret $\mathbf{w}$ as dynamic variables, proposing, as a possible implementation, that such non-linear functions are computed by two neural networks $W_{\Theta^{w}}$, and $Y_{\Theta^{y}}$:
$\begin{cases}e&=(y^{targ}-y)\\ \tau_{z}\dot{\mathbf{z}}&=\phi\left(J\mathbf{z}+R\mathbf{x}\right)-\mathbf{z}\\ y&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{\mathbf{x}|\mathbf{w}\})\simeq Y(\mathbf{x},\mathbf{w})=\mathbf{x}\cdot\mathbf{w}\\ \tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{\mathbf{x}|e\})\simeq W(\mathbf{x},e)=\mathbf{x}\,e\end{cases}$ (8)
In particular, we used two GMRNNs that receive $\mathbf{x},\mathbf{w},e$ as inputs and provide the proper output thanks to a suited training of their readout weights ($\Theta^{y}$ and $\Theta^{w}$, following reservoir computing paradigm). Notice that now (8) represents an RNN dynamically performing the learning process.
Dynamical reinforcement learning
In the context of reinforcement learning, our framework outlines a systematic approach for modelling agents that can dynamically adapt their behavior across diverse environments. We assume the same state transition dynamics across all the environments $\dot{\mathbf{x}}=X(\mathbf{x},a)$, while the reward distribution $r\sim Prob_{\alpha}(r|\mathbf{x},a)$ depends on the current environment $\alpha\in\mathcal{A}$. Here $a(t)$ is an integer value indicating the action at time $t$ among $\mathrm{D}$ possible ones.
We start by defining a policy network, denoted as $\mathbf{\pi}=\mathrm{softmax}(\mathbf{y})$ which implements a policy mapping the agent state encoded by the feature vector $\mathbf{z}$, to a probability distribution over actions. We assume an agent that linearly recombines these features $\mathbf{y}=Y(\mathbf{z},\mathbf{w})=\mathbf{w}\cdot\mathbf{z}$. This assumption is made without loss of generality, since we could always employ a reservoir computer as an intermediate feature extraction layer, as described in the previous subsection. The agent learns by evaluating the policy gradient with respect to the weights $\mathbf{w}$. This leads to a dynamics similar to equation (8), where the definition of $e(t)$ is changed as follows:
$\begin{cases}\mathbf{e}&=r\left(\mathbf{\mathds{1}}_{a}-\mathbf{\pi}\right)\\ \mathbf{y}&=\mathbf{w}\cdot\mathbf{z}\\ \tau_{w}\,\dot{\mathbf{w}}&=\mathbf{e}\odot\mathbf{z}\end{cases}$ (9)
Here $\mathds{1}_{a(t)}$ represents the ’one-hot encoded’ action (as defined in [52, 53]) at time $t$. It is a $\mathrm{D}$-element vector where the $a(t)$-th element is one, and all other elements are zero. This formulation holds for a null discount factor [52, 53]), we refer to the Appendix section Reinforcement Learning and the Discount Factor for the description of the general case.
As before, we model that the policy is modulated by an auxiliary population of virtual weights $\mathbf{w}$, determined by the activity of an additional GMRNN $W_{\Theta_{y}}$. This auxiliary network adjusts its internal activity based on the rewards received, effectively implementing policy gradient updates of the virtual weights and thereby modulating the agent behavior in real-time. Then, the above dynamics can be approximated by two networks, one for estimating the gradients, and one for the scalar product.
$\begin{cases}\tau_{w}\,\dot{\mathbf{w}}&={\color[rgb]{0,0.66796875,0}W_{\Theta^{w}}}(\{\mathbf{z}|r,a,\mathbf{\pi}\})\simeq W(\mathbf{z},r,a,\mathbf{\pi})=r\left(\mathbf{\mathds{1}}_{a}-\mathbf{\pi}(\mathbf{y})\right)\odot\mathbf{z}\\ \mathbf{y}&={\color[rgb]{1,.5,0}Y_{\Theta^{y}}}(\{\mathbf{z}|\mathbf{w}\})\simeq Y(\mathbf{z},\mathbf{w})=\mathbf{x}\cdot\mathbf{w}\end{cases}$ (10)
Acknowledgments